OpenStudy (fionacndg):

A floor plan is given for the first floor of a new house. One inch represents 15 feet. If the living room is three fourths in. by two thirds in. on the floor plan, what are the actual dimensions?

OpenStudy (anonymous):

11.25 ft. by 10 ft.

OpenStudy (anonymous):

15/4 gives us a number for how many feet one forth of an inch is. 15/3 gives us a number for how many feet one third of an inch is. 15/4=3.75 3.75*3=11.25 15/3=3 3*2=10

OpenStudy (fionacndg):

thank you so much! I understand how to find the answer now! @OrangeMaster

OpenStudy (anonymous):

No problem. :) Glad you know how to do it now.